# The evolution of the galaxy stellar mass function and star formation rates in the COLIBRE simulations from redshift 17 to 0

### v1

@@ 4 Conclusions
In this work, we studied the evolution of stellar masses and star formation rates in new colibre cosmological hydrodynamical simulations of galaxy formation (Schaye et al., 2025; Chaikin et al., 2025). We analyzed a suite of fiducial colibre simulations covering a range of cosmological volumes and resolutions: $(400~\mathrm{cMpc})^{3}$ and $(200~\mathrm{cMpc})^{3}$ at m7 resolution with a gas particle mass of $m_{\rm gas}=1.47\times 10^{7}~\mathrm{M_{\odot}}$; $(200~\mathrm{cMpc})^{3}$ at m6 resolution with $m_{\rm gas}=1.8\times 10^{6}~\mathrm{M_{\odot}}$; and $(50~\mathrm{cMpc})^{3}$ and $(25~\mathrm{cMpc})^{3}$ at m5 resolution with $m_{\rm gas}=2.3\times 10^{5}~\mathrm{M_{\odot}}$. Our main focus was the evolution of the galaxy stellar mass function (GSMF), which we predicted over $0<z<17$. Additionally, we investigated the evolution of the stellar-to-halo mass relation (SHMR), star-forming main sequence (SFMS), galaxy quenched fraction, cosmic star formation rate density (CSFRD), cosmic stellar mass density (CSMD), and the number density of massive quiescent galaxies at high redshifts. To facilitate a fair comparison with observations, where inferred stellar masses have non-negligible errors, the GSMF, SFMS, galaxy quenched fraction, and number density of massive quiescent galaxies were computed using stellar masses of simulated galaxies that include a lognormal scatter with a standard deviation given by equation (1). Finally, we studied the differences between two AGN feedback implementations that are available in colibre: the fiducial model with thermal energy injections and a hybrid model combining thermal and kinetic jet feedback. Our key findings are as follows:
  • The colibre GSMF is consistent with the observationally inferred data over the entire redshift range where the observations are available ($0<z<12$), while demonstrating excellent convergence with resolution over a factor of $64$ in gas particle mass (Figs. 1 and 2). The largest systematic deviations from the observations, $\approx 0.3$ dex, occur at $z\approx 2-4$. At much higher redshifts ($12<z<17$), which lack observational measurements, colibre predicts a smoothly evolving GSMF, with the normalization changing by $\approx 2$ dex between $z=12$ and $z=17$ (Fig. 2). At the highest redshift probed by the simulations, $z=17$, the combined GSMF predicted by the colibre m5 and m6 resolutions still spans more than 2 dex in stellar mass.
@@
  • The CSFRD reproduces observationally inferred values reasonably well for $0<z<18$ (Fig. 5). In particular, at $8<z<18$, colibre predictions for the CSFRD from subhaloes with $M_{*}\geq 10^{7}~\mathrm{M_{\odot}}$ are consistent with JWST UV measurements reported by Harikane et al. (2023), Donnan et al. (2024), and Weibel et al. (2025a). At $z\approx 2$, colibre captures some observational trends while underestimating others by 0.20.50.2-0.5 dex, reflecting the well-known tension between intrinsic and some observed SFRs during this epoch. By $z=0$, the predictions are again consistent with observations, yielding a total CSFRD of $\approx 0.012~\mathrm{M_{\odot}\,yr^{-1}\,cMpc^{-3}}$.
@@
  • Similarly to the CSFRD, colibre reproduces the observed SFMS reasonably well at $0<z<1$ and $z>4$, but underpredicts the data at intermediate redshifts (Fig. 8). Specifically, between $1<z<4$, colibre predicts an SFMS that is lower by $\approx 0.2-0.6$ dex compared to the values reported by Thorne et al. (2021) and Popesso et al. (2023), although it agrees much better with the measurements by Leja et al. (2022). At $z<5$, higher-resolution simulations yield an SFMS with a slightly lower normalization, leading to somewhat worse agreement with observations, whereas at $z>5$, the convergence with resolution is excellent.
@@
  • colibre is consistent with recent measurements by JWST of the number density of massive quiescent galaxies at $2<z<7$ (Fig. 11). At $z\approx 7$, colibre predicts a comoving number density of galaxies with $M_{*}>10^{10}~\mathrm{M_{\odot}}$ and $\mathrm{sSFR}<10^{-10}~\mathrm{yr^{-1}}$ of $\sim 10^{-6.5}~\mathrm{cMpc^{-3}}$, which rises to $\sim 10^{-3.5}~\mathrm{cMpc^{-3}}$ by $z\approx 2$. Moreover, in the largest colibre volume ($400^{3}~\mathrm{cMpc}^{3}$), quiescent galaxies with $M_{*}>10^{11}~\mathrm{M_{\odot}}$ are found out to $z=7.5$, with a number density of $\sim 10^{-8}~\mathrm{cMpc^{-3}}$ at $z\approx 7.5$ and $\sim 10^{-6}~\mathrm{cMpc^{-3}}$ at $z\approx 5$. At a fixed redshift and stellar mass bin, the m6 model predicts a somewhat lower number density than m7, with differences ranging between $\approx 0.2$ and $1$ dex, while the volume of the largest available m5 simulation ($50^{3}~\mathrm{cMpc}^{3}$) is too small to make meaningful predictions at these high redshifts.
@@
Follow-up work will investigate early quenching mechanisms of massive galaxies in colibre and compare low- and high-redshift luminosity functions predicted by the simulations with observational data, to provide further theoretical insights into the evolution of galaxy stellar mass and star formation rates in the observable Universe.
@@ Acknowledgements
We thank Yannick M. Bahé, Simon Driver, Victor J. Forouhar Moreno, Anna de Graaff, Claudia del P. Lagos, and John Weaver for their useful comments. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. This project has received funding from the Netherlands Organization for Scientific Research (NWO) through research programme Athena 184.034.002. ABL acknowledges support by the Italian Ministry for Universities (MUR) program ‘Dipartimenti di Eccellenza 2023-2027’ within the Centro Bicocca di Cosmologia Quantitativa (BiCoQ), and support by UNIMIB’s Fondo Di Ateneo Quota Competitiva (project 2024-ATEQC-0050). CSF acknowledge support from European Research Council (ERC) Advanced Grant DMIDAS (GA 786910). SP acknowledges support from the Austrian Science Fund (FWF) through project V 982-N. Data analysis was carried out using swiftsimio (Borrow & Borrisov, 2020; Borrow & Kelly, 2021).
@@ Appendix A Convergence with resolution and volume
Figure 14: The impact of numerical resolution (left) and cosmological volume (right) on the evolution of the cosmic stellar mass density (CSMD). The left-hand panels compare the m5, m6, and m7 resolution simulations in the ($25~\mathrm{cMpc}$)3 volume, while the right-hand panels compare ($25~\mathrm{cMpc}$)3, ($50~\mathrm{cMpc}$)3, ($100~\mathrm{cMpc}$)3, and ($200~\mathrm{cMpc}$)3 volumes at m6 resolution. The layout of the panels is the same as in Fig. 4. The CSMD exhibits excellent convergence with resolutions (cosmological volume) at $z<5$ ($z<4$), with relative differences remaining below $\approx 20$ per cent.
In this section, we test the convergence of the colibre simulations with numerical resolution and with cosmological volume. Because some subgrid parameter values in the colibre models differ slightly between resolutions (see table 1 in Schaye et al. 2025, for a summary of these differences), convergence with resolution should be understood as ‘weak convergence’, following the terminology of Schaye et al. (2015).
@@ A.1 Convergence of the cosmic stellar mass density
Figure 15: As Fig. 11, but shows the comoving number density of massive quiescent galaxies predicted by the colibre m7 simulations in different cosmological volumes: ($\mathrm{50~cMpc}$)3, ($\mathrm{100~cMpc}$)3 ($\mathrm{200~cMpc}$)3, and ($\mathrm{400~cMpc}$)3, indicated by different shades of red. For reference, the thin horizontal dotted lines mark the minimum non-zero number density of quiescent galaxies that each volume can resolve. While predictions from volumes of $(100~\mathrm{cMpc})^{3}$ and larger converge at $z\lesssim 4.5$, the $(50~\mathrm{cMpc})^{3}$ volume is too small to provide meaningful predictions at $z\gtrsim 3.5$.
The CSMD exhibits excellent convergence with resolution at $z\lesssim 5$, with relative differences between adjacent resolutions within $\approx 20$ per cent. This excellent convergence holds both for the total CSMD, computed using all stellar particles in the simulation (solid curves), and for the CSMD contributed only by subhaloes with stellar masses $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ (dash-dotted curves). At higher redshifts ($z\gtrsim 6$), the total CSMDs begin to diverge, with higher-resolution simulations predicting systematically higher values of the CSMD. This trend is expected, as the total CSMD at these high redshifts is dominated by SFRs of the least massive subhaloes (see Fig. 6), whose abundance is sensitive to numerical resolution.
The CSMD also shows excellent convergence with cosmological volume at fixed resolution for $z\lesssim 4$, which holds for both the total CSMD and that contributed only by $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ subhaloes. At higher redshifts, convergence remains good for the total CSMD up to $z\approx 14$, with relative differences between simulations in different volumes remaining within $\approx 30$ per cent. This behaviour is expected, as at high redshifts the total CSMD is dominated by the least massive objects ($M_{*}\lesssim 10^{6.5}~\mathrm{M_{\odot}}$; see Fig. 23), which are abundant in all four volumes. In contrast, the CSMD from $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ subhaloes depends sensitively on cosmological volume above $z\approx 4$. On average, larger volumes yield higher CSMDs at high redshifts, as they are more likely to host $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ galaxies. One exception to this trend is the ($100~\mathrm{cMpc}$)3 volume, which predicts a slightly higher CSMD than that in the larger ($200~\mathrm{cMpc}$)3 volume due to a rare overdensity in its ICs that forms a massive protocluster by $z\sim 5$, atypical for a ($100~\mathrm{cMpc}$)3 volume.
Figure 16: As in Fig. 10, but with all simulations run in the same cosmological volume of ($25~\mathrm{cMpc}$)3. Using the same cosmological volume at all colibre resolutions improves the resolution convergence of the quenched fraction for satellite galaxies, compared to Fig. 10, where the volume of the m5 simulation is much smaller than in the m6 and m7 counterparts.
@@ A.2 Convergence of the number density of massive quiescent galaxies
While predictions from volumes of $(100~\mathrm{cMpc})^{3}$ or larger converge at $z\lesssim 4.5$, the $(50~\mathrm{cMpc})^{3}$ volume is too small to yield meaningful predictions above $z\approx 3$. In particular, its minimum resolvable number density of $\sim 10^{-5}~\mathrm{cMpc}^{-3}$ makes it unsuitable for comparison with high-redshift ($z\gtrsim 4$) observations, which typically report number densities comparable to or lower than $\sim 10^{-5}~\mathrm{cMpc}^{-3}$.
@@ A.3 Convergence of the galaxy quenched fraction
Fig. 16 shows the $z=0$ quenched fraction for all, central, and satellite galaxies predicted by the simulations at m7, m6, and m5 resolutions, all using the ($25~\mathrm{cMpc}$)3 cosmological volume. We find that the resolution convergence of the quenched fraction for satellite galaxies between the m5 and the two higher resolutions improves significantly when using the same cosmological volume, compared to Fig. 10, which shows the quenched fraction in the simulations at different resolutions in different volumes.
@@ D.3 The effect of Eddington bias on the number density of massive quiescent galaxies
Fig. 21 shows the effect of Eddington bias on the evolution of the comoving number density of massive quiescent galaxies ($n_{\rm q}$) in the L200m6 simulation. The figure layout matches Fig. 11, including the observational data (black symbols). As in Fig. 11, different panels correspond to different stellar mass bins for quiescent galaxies contributing to $n_{\rm q}$: $M_{*}>10^{10}~\mathrm{M_{\odot}}$ (top), $M_{*}>10^{10.5}~\mathrm{M_{\odot}}$ (middle), and $M_{*}>10^{11}~\mathrm{M_{\odot}}$ (bottom). In the simulation, quiescent galaxies satisfy $\mathrm{sSFR}<10^{-10}~\mathrm{yr}^{-1}$. Note that while curves in Fig. 21 correspond to fixed $\sigma_{\rm random}$ values (as indicated in the legend), Fig. 11 uses $\sigma_{\rm random}$ from equation (1), which depends on redshift but saturates at $0.3$ dex for $z\geq 2$. Adding lognormal scatter affects both galaxy stellar masses and their sSFRs, since sSFR is defined relative to stellar mass.

### v2

@@ 4 Conclusions
In this work, we have studied the evolution of stellar masses and star formation rates in new colibre cosmological hydrodynamical simulations of galaxy formation (Schaye et al., 2026; Chaikin et al., 2026). We analyzed a suite of fiducial colibre simulations covering a range of cosmological volumes and resolutions: $(400~\mathrm{cMpc})^{3}$ and $(200~\mathrm{cMpc})^{3}$ at m7 resolution with a gas particle mass of $m_{\rm gas}=1.47\times 10^{7}~\mathrm{M_{\odot}}$; $(200~\mathrm{cMpc})^{3}$ at m6 resolution with $m_{\rm gas}=1.8\times 10^{6}~\mathrm{M_{\odot}}$; and $(100~\mathrm{cMpc})^{3}$, $(50~\mathrm{cMpc})^{3}$, and $(25~\mathrm{cMpc})^{3}$ at m5 resolution with $m_{\rm gas}=2.3\times 10^{5}~\mathrm{M_{\odot}}$. Our main focus was the evolution of the galaxy stellar mass function (GSMF), which we predicted over $0<z<17$. Additionally, we investigated the evolution of the stellar-to-halo mass relation (SHMR), star-forming main sequence (SFMS), galaxy quenched fraction, cosmic star formation rate density (CSFRD), cosmic stellar mass density (CSMD), and the number density of massive quiescent galaxies at high redshift. To facilitate a fair comparison with observations, where inferred stellar masses have non-negligible errors, the GSMF, SFMS, galaxy quenched fraction, and number density of massive quiescent galaxies were computed using stellar masses of simulated galaxies that include a lognormal scatter with a standard deviation given by equation (1). Finally, we studied the differences between two AGN feedback implementations that are available in colibre: the fiducial model with thermal energy injections and a hybrid model combining thermal and kinetic jet feedback. Our key findings are as follows:
  • The colibre GSMF is consistent with the observationally inferred data over the entire redshift range where the observations are available ($0<z<12$), while demonstrating very good convergence with resolution over a factor of $64$ in gas particle mass (Figs. 1 and 2). The largest systematic deviations from the observations, $\approx 0.3$ dex, occur at $z\approx 2-4$. At much higher redshifts ($12<z<17$), which lack observational measurements, colibre predicts a smoothly evolving GSMF, with the normalization changing by $\approx 2$ dex between $z=12$ and $z=17$ (Fig. 2). At the highest redshift probed by the simulations, $z=17$, the combined GSMF predicted by the colibre m5 and m6 resolutions still spans more than 2 dex in stellar mass.
@@
  • The CSFRD reproduces observationally inferred values reasonably well for $0<z<18$ (Fig. 5). In particular, at $8<z<18$, colibre predictions for the CSFRD from subhaloes with $M_{*}\geq 10^{7}~\mathrm{M_{\odot}}$ are consistent with JWST UV measurements reported by Harikane et al. (2023), Donnan et al. (2024), and Weibel et al. (2025a). At $z\approx 2$, colibre captures some observational trends while underestimating others by 0.20.50.2-0.5 dex, reflecting the well-known tension between intrinsic and some observed SFRs during this epoch. By $z=0$, the predictions are again consistent with observations, yielding a total CSFRD of $\approx 0.012-0.013~\mathrm{M_{\odot}\,yr^{-1}\,cMpc^{-3}}$.
@@
  • Similarly to the CSFRD, colibre reproduces the observed SFMS reasonably well at $0<z<1$ and $z>4$, but underpredicts the data at intermediate redshifts (Fig. 8). Specifically, between $1<z<4$, colibre predicts an SFMS that is lower by $\approx 0.2-0.6$ dex compared to the values reported by Thorne et al. (2021) and Popesso et al. (2023), although it agrees much better with the measurements by Leja et al. (2022). At $z<5$, higher-resolution simulations yield an SFMS with a slightly yet systematically lower normalization, leading to somewhat worse agreement with observations, whereas at higher redshifts, the convergence with resolution is generally good.
@@
  • colibre is consistent with recent measurements by JWST of the number density of massive quiescent galaxies at $2<z<7$ (Fig. 11). At $z\approx 7$, colibre predicts a comoving number density of galaxies with $M_{*}>10^{10}~\mathrm{M_{\odot}}$ and $\mathrm{sSFR}<10^{-10}~\mathrm{yr^{-1}}$ of $\sim 10^{-6.5}~\mathrm{cMpc^{-3}}$, which rises to $\sim 10^{-3.5}~\mathrm{cMpc^{-3}}$ by $z\approx 2$. Moreover, in the largest colibre volume ($400^{3}~\mathrm{cMpc}^{3}$), quiescent galaxies with $M_{*}>10^{11}~\mathrm{M_{\odot}}$ are found out to $z=7.5$, with a number density of $\sim 10^{-8}~\mathrm{cMpc}^{-3}$ at $z\approx 7.5$ and $\sim 10^{-6}~\mathrm{cMpc}^{-3}$ at $z\approx 5$. Predictions from the largest available m5 simulation ($100^{3}~\mathrm{cMpc}^{3}$) are available only over a very narrow redshift range ($5<z<6$), owing to the relatively small cosmological volume and because the simulation has thus far progressed only down to $z=4.6$. At a fixed redshift and stellar mass bin, the m6 model predicts a somewhat lower number density than the m7 model, with differences ranging between $\approx 0.2$ and $1$ dex.
@@
We also note that, given the likely presence of systematic uncertainties in the observational analyses, we cannot exclude the possibility that some of the agreement between colibre and the data found in this work is fortuitous. In future work, we will perform forward modelling to compare colibre predictions with data in observer space, providing an important consistency check. Further follow-up work will investigate early quenching mechanisms of massive galaxies in colibre and compare low- and high-redshift luminosity functions predicted by the simulations with observational data, to provide further theoretical insights into the evolution of galaxy stellar mass and star formation rates in the observable Universe.
@@ Acknowledgements
We thank Yannick M. Bahé, Simon Driver, Victor J. Forouhar Moreno, Anna de Graaff, Claudia del P. Lagos, Christopher Lovell, Kai Wang, John Weaver, and the anonymous referee for their useful comments. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. This project has received funding from the Netherlands Organization for Scientific Research (NWO) through research programme Athena 184.034.002. ABL acknowledges support by the Italian Ministry for Universities (MUR) program ‘Dipartimenti di Eccellenza 2023-2027’ within the Centro Bicocca di Cosmologia Quantitativa (BiCoQ), and support by UNIMIB’s Fondo Di Ateneo Quota Competitiva (project 2024-ATEQC-0050). CSF acknowledge support from European Research Council (ERC) Advanced Grant DMIDAS (GA 786910). SP acknowledges support from the Austrian Science Fund (FWF) through project V 982-N. Data analysis was carried out using swiftsimio (Borrow & Borrisov, 2020; Borrow & Kelly, 2021).
@@ Appendix A Convergence with resolution and volume
Figure 14: The impact of numerical resolution (left) and cosmological volume (right) on the evolution of the cosmic stellar mass density (CSMD). The left-hand panels compare the m5, m6, and m7 resolution simulations in the ($25~\mathrm{cMpc}$)3 volume, while the right-hand panels compare ($25~\mathrm{cMpc}$)3, ($50~\mathrm{cMpc}$)3, ($100~\mathrm{cMpc}$)3, and ($200~\mathrm{cMpc}$)3 volumes at m6 resolution. The layout of the panels is the same as in Fig. 4. The CSMD exhibits very good convergence with resolution (cosmological volume) at $z<5$ ($z<4$), with relative differences remaining below $\approx 20$ per cent.
In this section, we test the convergence of the colibre simulations with numerical resolution and with cosmological volume. Because some subgrid parameter values in the colibre models differ slightly between resolutions (see table 1 in Schaye et al. 2026, for a summary of these differences), convergence with resolution should be understood as ‘weak convergence’, following the terminology of Schaye et al. (2015).
We emphasize that in colibre, the strength of SN and AGN feedback was optimized to reproduce the observed $z=0$ GSMF and the size – stellar mass relation in the stellar mass range $10^{9}<M_{*}/\mathrm{M}_{\odot}<10^{11.3}$ (see Chaikin et al., 2026, for details). Furthermore, the coupling efficiency of AGN feedback was tuned to match the observed $z=0$ BH mass – stellar mass relation of massive galaxies. The calibration was performed independently at each resolution, and no constraints at higher redshift were applied. Consequently, the colibre predictions at higher redshifts or outside the mass range used for calibration are not guaranteed to converge between different resolution levels, particularly for statistics other than those used in the calibration.
@@ A.1 Convergence of the cosmic stellar mass density
Figure 15: As Fig. 11, but shows the comoving number density of massive quiescent galaxies predicted by the colibre m7 simulations in different cosmological volumes: ($\mathrm{50~cMpc}$)3, ($\mathrm{100~cMpc}$)3 ($\mathrm{200~cMpc}$)3, and ($\mathrm{400~cMpc}$)3, indicated by different shades of red. For reference, the thin horizontal dotted lines mark the minimum non-zero number density of quiescent galaxies that each volume can resolve. While predictions from volumes of $(100~\mathrm{cMpc})^{3}$ and larger converge at $z\lesssim 5$, the $(50~\mathrm{cMpc})^{3}$ volume is too small to provide meaningful predictions at $z\gtrsim 3.5$.
The CSMD exhibits very good convergence with resolution at $z\lesssim 5$, with relative differences between adjacent resolutions within $\approx 20$ per cent. This convergence holds both for the total CSMD, computed using all stellar particles in the simulation (solid curves), and for the CSMD contributed only by subhaloes with stellar masses $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ (dash-dotted curves). At higher redshifts ($z\gtrsim 6$), the total CSMDs begin to diverge, with higher-resolution simulations predicting systematically higher values of the CSMD. This trend is expected, as the total CSMD at these high redshifts is dominated by SFRs of the least massive subhaloes (see Fig. 6), whose abundance is sensitive to numerical resolution.
The CSMD also shows very good convergence with cosmological volume at fixed resolution for $z\lesssim 4$, which holds for both the total CSMD and that contributed only by $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ subhaloes. At higher redshifts, convergence remains good for the total CSMD up to $z\approx 14$, with relative differences between simulations in different volumes remaining within $\approx 30$ per cent. This behaviour is expected, as at high redshifts the total CSMD is dominated by the least massive objects ($M_{*}\lesssim 10^{6.5}~\mathrm{M_{\odot}}$; see Fig. 23), which are abundant in all four volumes. In contrast, the CSMD from $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ subhaloes depends sensitively on cosmological volume above $z\approx 4$. On average, larger volumes yield higher CSMDs at high redshifts, as they are more likely to host $M_{*}\geq 10^{8}~\mathrm{M_{\odot}}$ galaxies. One exception to this trend is the ($100~\mathrm{cMpc}$)3 volume, which predicts a slightly higher CSMD than that in the larger ($200~\mathrm{cMpc}$)3 volume due to a rare overdensity in its ICs that forms a massive protocluster by $z\sim 5$, atypical for a ($100~\mathrm{cMpc}$)3 volume.
Figure 16: As in Fig. 10, but with all simulations using the same cosmological volume of ($25~\mathrm{cMpc}$)3. Adopting the same cosmological volume at all colibre resolutions improves the convergence of the quenched fraction for satellite galaxies, compared to Fig. 10, where the m5 simulation uses a much smaller volume than the m6 and m7 counterparts.
@@ A.2 Convergence of the number density of massive quiescent galaxies
While predictions from volumes of $(100~\mathrm{cMpc})^{3}$ or larger converge at $z\lesssim 5$, the $(50~\mathrm{cMpc})^{3}$ volume is too small to yield meaningful predictions above $z\approx 3.5$. In particular, its minimum resolvable number density of $\sim 10^{-5}~\mathrm{cMpc}^{-3}$ makes it unsuitable for comparison with high-redshift ($z\gtrsim 4$) observations, which typically report number densities comparable to or below this level. The $(100~\mathrm{cMpc})^{3}$ volume extends the predictions of non-zero number densities out to $z\approx 6.5$ for the adopted sSFR and $M_{*}$ cuts, with a minimum resolved number density of $10^{-6}~\mathrm{cMpc}^{-3}$; however, its predictions at $z\gtrsim 4$ are subject to small-number statistics.
@@ A.3 Convergence of the galaxy quenched fraction
Fig. 16 shows the $z=0$ quenched fraction for all, central, and satellite galaxies predicted by the simulations at m7, m6, and m5 resolutions, all using the $(25~\mathrm{cMpc})^{3}$ cosmological volume. We find that the convergence of the satellite quenched fraction between m5 and the two lower resolutions improves significantly when the same cosmological volume is used, compared to Fig. 10, which shows the quenched fraction in the simulations at different resolutions in different volumes.
@@ D.3 The effect of Eddington bias on the number density of massive quiescent galaxies
Fig. 21 shows the effect of Eddington bias on the evolution of the comoving number density of massive quiescent galaxies ($n_{\rm q}$) in the L200m6 simulation. The figure layout matches Fig. 11, including the observational data (black symbols). As in Fig. 11, different panels correspond to different stellar mass bins for quiescent galaxies contributing to $n_{\rm q}$: $M_{*}>10^{10}~\mathrm{M_{\odot}}$ (top), $M_{*}>10^{10.5}~\mathrm{M_{\odot}}$ (middle), and $M_{*}>10^{11}~\mathrm{M_{\odot}}$ (bottom). In the simulation, quiescent galaxies satisfy $\mathrm{sSFR}<10^{-10}~\mathrm{yr}^{-1}$. Note that while curves in Fig. 21 correspond to fixed $\sigma_{\rm random}$ values (as indicated in the legend), Fig. 11 uses $\sigma_{\rm random}$ from equation (1), which depends on redshift but saturates at $0.3$ dex for $z\geq 2$. Adding lognormal scatter affects both galaxy stellar masses and their sSFRs, as stellar mass appears in the definition of sSFR.